Operation Summary for May 11, 2011 - May 13, 2011 |
|
Wednesday May 11 The day shift began with Operations monitoring the Tevatron (TEV) store 8725 (144.01E30) and the Recycler stash (145.0E10), with the Antiproton Source (Pbar) stacking (6.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, T-978 ended their present run in the MTest line at 11:57 AM. TEV personnel began end-of-store (EOS) studies at 2 PM. Operations reported that a power glitch took down various RF stations and systems in the TEV, Pbar, and Recycler, but the store survived. The affected systems were largely recovered by 3:55 PM. During the evening shift, EOS studies concluded and operators terminated store 8725 at 6:04 PM. Operators prepared D0 and CDF for accesses; the experiments came out of access at 8:36 PM. Operations turned the TEV on, completed a check-out, and put the TEV into HEP shot setup. Operations established store 8728 at 11:02 PM with an average initial luminosity (AIL) of 361.7E30. Thursday May 12 The midnight shift began with Operations monitoring store 8728 (263.32E30) and the stash (70.9E10), with Pbar stacking (18.6E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 3:49 AM, a lightning strike resulted in the loss of 120E10 of antiprotons from the stash. The glitch also caused Booster RF trouble and caused the Tevatron Electron Lens (TEL1) to trip off – everything recovered quickly. At 6:10 AM, a MI quadrupole magnet power supply (I:QF12) tripped off on a ground fault indication; this kept beam from Pbar while operators and EE support personnel investigated. During the day shift, Operations terminated store 8728 at 10:08 AM. Operators prepared the TEV, MI, Pbar, Booster, and Linac for accesses at 10:30 AM. Operators and MI personnel accessed F- Sectors and reported finding the I:QF12 ground fault on the bus.
During the evening shift, Operations reestablished beam to MI at 5:09 PM, but a Pbar target change kept stacking off until 6:45 PM and controls issues kept full intensity beam from NuMI until 7:35 PM. The TEV accesses were completed and Operations had current back on the TEV bus by 6:43 PM. Operators conducted a TEV checkout and then put the TEV into shot setup. Friday May 13 The midnight shift began with the TEV in shot setup, with Operations monitoring the stash (117.5E10), with Pbar stacking (23.9E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 8733 at 1:26 AM with an AIL of 123.164E30. Operations established beam to the new Meson FTBF T-979 experiment at 7:59 AM. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 05/13/2011 email Fermilab |
| Security, Privacy, Legal | |